# Can Big Black Holes Merge with the Smallest Black Holes?

### v1

@@ Abstract
software: Bilby (Ashton et al., 2019), gwpopulation (Talbot et al., 2019), gwpopulation_pipe (Talbot, 2021), dynesty (Speagle, 2020), NumPy (Harris et al., 2020), matplotlib (Hunter, 2007), corner (Foreman-Mackey, 2016).
@@ 1 Introduction
GW observations of merging binary black holes from LIGO (Aasi et al., 2015), Virgo (Acernese et al., 2015), and KAGRA (Akutsu et al., 2021) allow us to probe the evolution of BBHs (Mapelli, 2020; Mandel & Farmer, 2022; Abbott et al., 2023a). The observed population likely contains contributions from multiple formation channels (Zevin et al., 2021; Cheng et al., 2023; Li et al., 2024; Colloms et al., 2025a), but the details of the astrophysical evolution of these systems are uncertain (Mandel & Broekgaarden, 2022; Belczynski et al., 2022; Mandel & Farmer, 2022). Measuring the underlying distribution of BBH masses (primary mass $m_{1}$ and secondary mass $m_{2}$), mass ratios ($q=m_{2}/m_{1}\leq 1$), spins, redshifts, and the correlations between these parameters from current gravitational-wave (GW) observations can inform us of formation mechanisms (Mapelli, 2021; Abbott et al., 2023a; Callister, 2024).
@@
BBH formation processes are expected to leave signatures in the distribution of component masses and mass ratios of merging binaries. The mass ratios of binary stars during their evolution depends on the initial fragmentation, accretion during star formation and evolution, and supernovae mechanisms (Bate & Bonnell, 1997; Belczynski et al., 2012; Moe & Di Stefano, 2017; Neijssel et al., 2019; Zevin et al., 2020; Broekgaarden et al., 2022; Briel et al., 2022). Population-synthesis simulations indicate that field binaries that undergo a common-envelope stage can produce BBHs with mass-ratio distributions with support down to $q\sim 0.2$, though this depends on the treatment of the common-envelope process (Zevin et al., 2020; Olejak et al., 2021). Simulations of binaries that undergo stable mass transfer predict this channel cannot produce such unequal mass binaries, but can produce preferentially more massive black holes than common envelope (van Son et al., 2022). In contrast, BBHs that merge due to dynamical encounters may have more unequal masses and higher masses than isolated formation channels, especially if they are hierarchical mergers (Rodriguez et al., 2019; Sedda et al., 2023, 2020; Bruel et al., 2024). The von Zeipel–Lidov–Kozai effect for field triples can produce BBHs with mass ratios down to $0.3$ (Martinez et al., 2022). Similarly, active galactic nuclei can produce BBHs with more unequal masses than isolated binaries (Yang et al., 2019; Tagawa et al., 2020; Secunda et al., 2020). High-mass BBHs with unequal-masses could therefore be a signature of formation via dynamical encounters or more exotic processes.
@@
Investigations of the mass distributions of black holes have found structure in the primary masses, secondary masses, and the mass ratios of the population. Fishbach et al. (2018) found that black holes do not randomly merge with each other regardless of mass; rather BBHs prefer equal masses. This is seen in the inferred mass-ratio distribution, which peaks at equal masses when considering a power-law model (Abbott et al., 2021, 2023a). Despite this preference for equal masses, secondary masses have been shown to follow a different distribution than primary masses, indicating that primary and secondary black holes undergo different astrophysical processes (Farah et al., 2024). Works considering more flexible models for the mass-ratio distribution find that the distribution peaks away from equal masses (Sadiq et al., 2024; Godfrey et al., 2023; Rinaldi et al., 2024; Callister & Farr, 2024a; Rinaldi et al., 2025). Furthermore, there is evidence that the mass-ratio distribution may change with primary mass (Sadiq et al., 2024) and that different features in the primary-mass distribution have different mass-ratio distributions (Li et al., 2022; Godfrey et al., 2023; Galaudage & Lamberts, 2025). These results indicate that there is information in the mass distributions that could be used to uncover the uncertain astrophysics of BBH formation (e.g., Colloms et al., 2025a). Exploring parametric model variations for BBH masses allows us to test assumptions about the relationship between component masses.
@@
Callister (2024) highlights that the correlations between parameters, in addition to the parameters’ marginal distributions, can be informative. When investigating correlations, it is necessary to check whether results are accurate descriptions of the population, or whether they are driven by the enforced model constraints. Adamcewicz & Thrane (2022) investigated the anti-correlation between mass ratio and effective inspiral spin. By fixing the shape of the marginal distributions, they avoid the case where the correlation results from an improved fit of the marginal distributions. The same can be true in the other direction: inferences of marginal distributions could be driven by assumed correlations (or the lack thereof) imposed by the chosen model. Potentially, individual source constraints that are not well accounted for by model correlations could influence the measured marginal distributions, especially if a correlation must be distorted to fit them. Updating population models so that outliers are well accounted for and correlations are appropriately incorporated helps us understand the entire population of BBHs.
@@
Modelling BBHs mass distributions requires a minimum black hole mass. This raises the question of the classification of merging objects with a mass that could correspond either to a black hole or a neutron star (Farah et al., 2022). It is unknown whether the maximum neutron star mass is the same as the maximum black hole mass. The maximum neutron star mass is uncertain, and depends on the neutron star equation of state (Rutherford et al., 2024; Koehn et al., 2025). Meanwhile, uncertainties in the minimum black hole mass come from the details of supernovae and core collapse (Fryer et al., 2012; Liu et al., 2021; Patton et al., 2022; Fryer et al., 2022). Following X-ray binary observations, it has been proposed that the maximum neutron star mass and the minimum black hole mass are not equal (Bailyn et al., 1998; Ozel et al., 2010; Farr et al., 2011; Siegel et al., 2023), and there is a mass gap with a dearth of black holes. Observations of objects with masses $\sim 2.5$$5~{M}_{\odot}$ indicate that at least some of the proposed gap is populated, though the rate of these objects is uncertain (Thompson et al., 2019; Abbott et al., 2020; Jayasinghe et al., 2021; Abbott et al., 2023b; Abac et al., 2024; Fishbach et al., 2025). Including low-mass black holes from GW observations can allow us to measure the minimum black hole mass (Ray et al., 2025).
@@
Observations of (potential) low-$m_{2}$ BBHs include GW190814, which has $q=0.112^{+0.008}_{-0.009}$ and $m_{2}=2.6^{+0.1}_{-0.1}{M}_{\odot}$ (Abbott et al., 2020, 2024). Understanding the nature of GW190814’s source as a BBH or a neutron star–black hole requires understanding if it is an outlier to the BBH population, and whether its outlier status is because of population misspecification. While Abbott et al. (2021) and Essick et al. (2022) find GW190814 to be an outlier, Farah et al. (2022) consider models where it is not by including the full population of neutron stars and black holes, and considering a partially full mass gap. The outlier status of observations thus depends on the choice of population model. We seek to understand how to improve our models and incorporate observations that are currently challenging to fit.
@@
Current models for the BBHs mass distribution assume that the minimum black hole mass is constant with the primary black hole mass. Given the preference for equal mass mergers and the structure in the distribution of BBH masses, this may not be the case. We investigate if there is evidence for an evolving minimum secondary mass with primary mass: can black holes with higher masses merge with the lowest-mass black holes? We introduce models for the mass-ratio distribution of BBHs specifying a minimum secondary black hole mass that evolves with the primary mass (Section 2). Using these models on observations from the third Gravitational-Wave Transient Catalog (GWTC-3.0; Abbott et al., 2023b) we infer the primary-mass and mass-ratio distributions (Section 3). Our conclusions for the inferred evolution of minimum secondary mass, and measured primary-mass and mass-ratio distributions are discussed in Section 4.
@@ 2 Methods
We estimate a set of hyperparameters $\Lambda$ that govern the shape of the mass, spin and redshift distributions of GW observations from third Gravitational-Wave Transient Catalog (GWTC-3.0) (Abbott et al., 2023b) using a hierarchical Bayesian inference (Mandel et al., 2019; Thrane & Talbot, 2019; Vitale et al., 2020). For each population distribution we choose a parametric model. We assume the distribution of the primary mass, spin and redshift distributions to be the default parametric models used in Abbott et al. (2023a), and introduce new models for the mass-ratio distribution as described in Section 2.1. Our inference framework is described in more detail in section Section 2.2 and the observations we use are described in Section 2.3.
@@ 2.1 Variable minimum-mass models
To investigate the correlation between the minimum mass and the primary mass, we model the mass-ratio distribution of BBHs with a minimum secondary mass that varies with the primary mass. We include this as a modification to a power-law model of mass ratio (Fishbach & Holz, 2017; Kovetz et al., 2017), such that the mass-ratio distribution ${\mathrm{p}}(q\mid m_{1})$ with a power-law index $\beta$ and a minimum secondary mass $m_{2}^{\mathrm{min}}(m_{1})$ follows
$\begin{split}{\mathrm{p}}(q\mid m_{1})&=\left(1+\beta\right)\left[1-\frac{m_{2}^{\mathrm{min}}(m_{1})}{m_{1}}\right]^{-(1+\beta)}q^{\beta}\\ &\propto q^{\beta}\qquad\qquad\left\{\frac{m_{2}^{\mathrm{min}}(m_{1})}{m_{1}}\leq q\leq 1\right\}.\end{split}$ (1)
This allows a varying lower bound on the secondary mass $m_{2}$ such that $m_{2}\geq m_{2}^{\mathrm{min}}(m_{1})$: the smallest possible secondary mass, and therefore the minimum mass ratio varies with the primary mass.
@@
This functional form automatically enforces that $m_{2}^{\mathrm{min}}(m_{\mathrm{min}})=m_{\mathrm{min}}$. We enforce additional constraints to ensure that the model remains physical. We require that $m_{2}^{\mathrm{min}}(m_{1})<m_{1}$ for all $m_{1}<m_{\mathrm{max}}$. Equivalently,
@@ 2.2 Statistical framework
where the mass-ratio distribution includes the Power Law, Increasing Parabola or Relaxed Parabola minimum secondary-mass model described above, and the other observables have distributions that follow the default parametric models used in Abbott et al. (2023a). We use the Power Law+Peak model for $p(m_{1})$ (Talbot & Thrane, 2018). This models the $m_{1}$ distribution as a two component mixture of a power-law of slope $\alpha$ and a Gaussian component. We choose the similar priors on the default model parameters as Abbott et al. (2023a), with some refinements. We use primary-mass power-law index $\alpha\sim\mathcal{U}(-4,12)$, mass-ratio power-law index $\beta\sim\mathcal{U}(-2,7)$, primary minimum mass $m_{\mathrm{min}}\sim\mathcal{U}(2,6)\,M_{\odot}$, and maximum mass $m_{\mathrm{max}}\sim\mathcal{U}(62,100)\,M_{\odot}$. Both spin magnitudes are modelled with an identical beta distribution, and the spin tilts a mixture of isotropic and normally distributed spin tilts. The redshift distribution $p(z)$ is modelled with a power law (Fishbach et al., 2018) of index $\kappa\sim\mathcal{U}(-2,8)$ .
To infer hyperposteriors on $\Lambda$ we use the population-inference package gwpopulation_pipe (Talbot, 2021) with the nested sampler dynesty (Speagle, 2020). The inference imposes a cut on the likelihood variance of 11 during sampling (Talbot & Golomb, 2023).
@@ 2.3 Gravitational-wave observations
We use the observations from third Gravitational-Wave Transient Catalog (GWTC-3.0; Abbott et al., 2023b) selected with a false alarm rate threshold of less than $1~\mathrm{yr}^{-1}$. This is identical to the set of observations used in Abbott et al. (2023a). We consider only signals with BBH sources and the ambiguous GW190814 (Abbott et al., 2020, 2024). As GW190814’s secondary object may be above the expected maximum mass for a neutron star, we compare the results with and without GW190814, and investigate its potential as an outlier to the BBH population. GW200210_092254 (Abbott et al., 2023b), which has a similar secondary mass and mass ratio to GW190814, is not included in our analysis as it does not satisfy the false alarm rate threshold.
We use a combined injection set of semi-analytic injections at the sensitivity of the first and second observing runs, and real injections at the third observing run sensitivity to account for selection effects (Abbott et al., 2023a). We consider the real injections with a false alarm rate of less than $1~\mathrm{yr}^{-1}$ detectable, and the semi-analytic injections we cut in network signal-to-noise ratio of $10$ (Abbott et al., 2018).

### v2

@@ Abstract
software: Bilby (2019Bilby), gwpopulation (2019gwpop), gwpopulation_pipe (gwpop_pipe), dynesty (Speagle_2020), NumPy (harris2020array), matplotlib (Hunter:2007), corner (corner).
@@ 1 Introduction
GW observations of merging binary black holes from LIGO (LIGOScientific:2014pky), Virgo (VIRGO:2014yos), and KAGRA (KAGRA:2020tym) allow us to probe the evolution of BBHs (Mapelli:2020vfa; Mandel:2018hfr; KAGRA:2021duu). The observed population likely contains contributions from multiple formation channels (zevinOneChannelRule2021; Cheng:2023ddt; Li:2023yyt; Colloms:2025hib), but the details of the astrophysical evolution of these systems are uncertain (Mandel:2021smh; Belczynski:2021zaz; Mandel:2018hfr). Measuring the underlying distribution of BBH masses (primary mass $m_{1}$ and secondary mass $m_{2}$), mass ratios ($q=m_{2}/m_{1}\leq 1$), spins, redshifts, and the correlations between these parameters from current gravitational-wave (GW) observations can inform us of formation mechanisms (Mapelli:2021taw; KAGRA:2021duu; Callister:2024cdx).
@@
BBH formation processes are expected to leave signatures in the distribution of component masses and mass ratios of merging binaries. The mass ratios of binary stars during their evolution depend on the initial fragmentation, accretion during star formation and evolution, and supernovae mechanisms (Bate8134774; Belczynski:2011bn; Moe_2017; neijsselEffectMetallicityspecificStar2019; Zevin:2020gma; Broekgaarden:2022nst; brielUnderstandingHighmassBinary2022). Population-synthesis simulations indicate that field binaries that undergo a common-envelope stage can produce BBHs with mass-ratio distributions with support down to $q\sim 0.2$, though this depends on the treatment of the common-envelope process (Zevin:2020gma; Olejak:2021fti). Simulations of binaries that undergo stable mass transfer predict this channel cannot produce such unequal-mass binaries, but can produce preferentially more massive black holes than common envelope (vanSon:2021zpk). In contrast, BBHs that merge due to dynamical encounters may have more unequal masses and higher masses than isolated formation channels, especially if they are hierarchical mergers (Rodriguez:2019huv; Sedda:2021vjh; Sedda:2020vwo; Bruel:2023bxl). The von Zeipel–Lidov–Kozai effect for field triples can produce BBHs with mass ratios down to $0.3$ (Martinez:2021tmr). Similarly, active galactic nuclei can produce BBHs with more unequal masses than isolated binaries (Yang:2019okq; Tagawa:2019osr; Secunda:2020mhd). High-mass BBHs with unequal masses could therefore be a signature of formation via dynamical encounters or more exotic processes.
@@
Investigations of the mass distributions of black holes have found structure in the primary masses, secondary masses, and the mass ratios of the population. Fishbach:2019bbm found that black holes do not randomly merge with each other regardless of mass; rather BBHs prefer equal masses. This is seen in the inferred mass-ratio distribution, which peaks at equal masses when considering a power-law model (LIGOScientific:2020kqk; KAGRA:2021duu). Farah:2023swu modelled the secondary-mass distribution distinctly from the primary-mass distribution, finding that while the distributions of primary and secondary masses are consistent, secondary masses may follow a different distribution than primary masses. Different distributions would indicate that primary and secondary black holes undergo different evolutionary processes. Works considering more flexible models for the mass-ratio distribution find that the distribution peaks away from equal masses (Sadiq:2023zee; godfreyCosmicCousinsIdentification2023; Rinaldi:2023bbd; Callister:2023tgi; Rinaldi:2025emt). Furthermore, there is evidence that the mass-ratio distribution may change with primary mass (Sadiq:2023zee) and that different features in the primary-mass distribution have different mass-ratio distributions (Li:2022jge; godfreyCosmicCousinsIdentification2023; Li:2024jzi; Galaudage:2024meo; Roy:2025ktr). These results indicate that there is information in the mass distributions that could be used to uncover the uncertain astrophysics of BBH formation (e.g., Colloms:2025hib). Exploring parametric model variations for BBH masses allows us to test assumptions about the relationship between component masses.
@@
Callister:2024cdx highlights that the correlations between parameters, in addition to the parameters’ marginal distributions, can be informative. When investigating correlations, it is necessary to check whether results are accurate descriptions of the population, or whether they are driven by the enforced model constraints. Adamcewicz:2022hce investigated the anti-correlation between mass ratio and effective inspiral spin. By fixing the shape of the marginal distributions, they avoid the case where the correlation results from an improved fit of the marginal distributions. The same can be true in the other direction: inferences of marginal distributions could be driven by assumed correlations (or the lack thereof) imposed by the chosen model. Potentially, individual source constraints that are not well accounted for by model correlations could influence the measured marginal distributions, especially if a correlation must be distorted to fit them. Updating population models so that outliers are well accounted for and correlations are appropriately incorporated helps us understand the entire population of BBHs.
@@
Modelling BBH mass distributions requires a minimum black hole mass. This raises the question of the classification of merging objects with a mass that could correspond either to a black hole or a neutron star (Farah:2021qom). It is unknown whether the maximum neutron star mass is the same as the minimum black hole mass. The maximum neutron star mass is uncertain, and depends on the neutron star equation of state (Rutherford:2024srk; Koehn:2024set). Meanwhile, uncertainties in the minimum black hole mass come from the details of supernovae and core collapse (Fryer:2011cx; Liu:2020uba; Patton:2021gwh; Fryer:2022lla). Following X-ray binary observations, it has been proposed that the maximum neutron star mass and the minimum black hole mass are not equal (Bailyn:1997xt; Ozel:2010su; Farr:2010tu; Siegel:2022gwc), and there is a mass gap with a dearth of black holes. Observations of objects with masses $\sim 2.5$$5~{M}_{\odot}$ indicate that at least some of the proposed gap is populated, though the rate of these objects is uncertain (Thompson:2018ycv; LIGOScientific:2020zkf; Jayasinghe:2021uqb; LIGOScientific:2021djp; LIGOScientific:2024elc; Fishbach:2025bjh). Including low-mass black holes from GW observations can allow us to measure the minimum black hole mass (Ray:2025aqr).
@@
Observations of (potential) low-$m_{2}$ BBHs include GW190814, which has $q=0.112^{+0.008}_{-0.009}$ and $m_{2}=2.6^{+0.1}_{-0.1}{M}_{\odot}$ (LIGOScientific:2020zkf; LIGOScientific:2021usb). Understanding the nature of GW190814’s source as a BBH or a neutron star–black hole requires understanding if it is an outlier to the BBH population, and whether its outlier status is because of population misspecification. While LIGOScientific:2020kqk and Essick:2021vlx find GW190814 to be an outlier, Farah:2021qom consider models where it is not by including the full population of neutron stars and black holes, and considering a partially full mass gap. The outlier status of observations thus depends on the choice of population model. We seek to understand how to improve our models and incorporate observations that are currently challenging to fit.
@@
Current models for the BBHs’ mass distribution assume that the minimum black hole mass is constant with the primary black hole mass. Given the preference for equal-mass mergers and the structure in the distribution of BBH masses, this may not be the case. We investigate if there is evidence for an evolving minimum secondary mass with primary mass: can black holes with higher masses merge with the lowest-mass black holes? We introduce models for the mass-ratio distribution of BBHs specifying a minimum secondary black hole mass that evolves with the primary mass (Section 2). Using these models on observations from the third Gravitational-Wave Transient Catalog (GWTC-3.0; LIGOScientific:2021djp) we infer the primary-mass and mass-ratio distributions (Section 3). Our conclusions for the inferred evolution of minimum secondary mass, and measured primary-mass and mass-ratio distributions are discussed in Section 4.
@@ 2 Methods
We estimate a set of hyperparameters $\Lambda$ that govern the shape of the mass, spin and redshift distributions of GW observations from third Gravitational-Wave Transient Catalog (GWTC-3.0) (LIGOScientific:2021djp) using a hierarchical Bayesian inference (mandelExtractingDistributionParameters2019; thraneIntroductionBayesianInference2019; vitaleInferringPropertiesPopulation2021). For each population distribution we choose a parametric model. We assume the distribution of the primary mass, spin and redshift distributions to be the default parametric models used in KAGRA:2021duu, and introduce new models for the mass-ratio distribution as described in Section 2.1. Our inference framework is described in more detail in section Section 2.2 and the observations we use are described in Section 2.3.
@@ 2.1 Variable minimum-mass models
To investigate the correlation between the minimum mass and the primary mass, we model the mass-ratio distribution of BBHs with a minimum secondary mass that varies with the primary mass. We include this as a modification to a power-law model of mass ratio (fishbachWhereAreLIGO2017; Kovetz:2016kpi), such that the mass-ratio distribution $p(q\mid m_{1})$ with a power-law index $\beta$ and a minimum secondary mass $m_{2}^{\mathrm{min}}(m_{1})$ follows
$\begin{split}p(q\mid m_{1})&=\left(1+\beta\right)\left[1-\frac{m_{2}^{\mathrm{min}}(m_{1})}{m_{1}}\right]^{-(1+\beta)}q^{\beta}\\ &\propto q^{\beta}\qquad\qquad\left\{\frac{m_{2}^{\mathrm{min}}(m_{1})}{m_{1}}\leq q\leq 1\right\}.\end{split}$ (1)
This allows a varying lower bound on the secondary mass $m_{2}$ such that $m_{2}\geq m_{2}^{\mathrm{min}}(m_{1})$, the smallest possible secondary mass, and therefore the minimum mass ratio varies with the primary mass.
@@
This functional form automatically enforces that $m_{2}^{\mathrm{min}}(m_{\mathrm{min}})=m_{\mathrm{min}}$. We apply additional constraints to ensure that the model remains physical. We require that $m_{2}^{\mathrm{min}}(m_{1})<m_{1}$ for all $m_{1}<m_{\mathrm{max}}$. Equivalently,
@@
Figure 1 shows graphical representations of each model for some fiducial hyperparameters.
Figure 1: Example evolution of $m_{2}^{\mathrm{min}}$ with $m_{1}$ for the Power Law, Increasing Parabola, and Relaxed Parabola model variations we consider, for some fiducial hyperparameters within our prior range. The shaded region indicates where $m_{1}>m_{2}^{\mathrm{min}}>0$.
@@ 2.2 Statistical framework
where the mass-ratio distribution includes the Power Law, Increasing Parabola or Relaxed Parabola minimum secondary-mass model described above, and the other observables have distributions that follow the default parametric models used in KAGRA:2021duu. We use the Power Law+Peak model for $p(m_{1})$ (Talbot:2018cva). This models the $m_{1}$ distribution as a two-component mixture of a power law of slope $\alpha$ and a Gaussian component. We choose similar priors on the default model parameters as KAGRA:2021duu, with some refinements. We use primary-mass power-law index $\alpha\sim\mathcal{U}(-4,12)$, mass-ratio power-law index $\beta\sim\mathcal{U}(-2,7)$, primary minimum mass $m_{\mathrm{min}}\sim\mathcal{U}(2,6)\,M_{\odot}$, and maximum mass $m_{\mathrm{max}}\sim\mathcal{U}(62,100)\,M_{\odot}$. Both spin magnitudes are modelled with an identical beta distribution (Wysocki:2018mpo), and the spin tilts a mixture of isotropic and normally distributed spin tilts (Talbot:2017yur). The redshift distribution $p(z)$ is modelled with a power law (Fishbach:2018edt) of index $\kappa\sim\mathcal{U}(-2,8)$ .
To infer hyperposteriors on $\Lambda$ we use the population-inference package gwpopulation_pipe (Talbot:2024yqw; gwpop_pipe) with the nested sampler dynesty (Speagle_2020). The inference imposes a cut on the likelihood variance of $1$ during sampling (Talbot:2023pex).
@@ 2.3 Gravitational-wave observations
We use the observations from GWTC-3.0 (LIGOScientific:2021djp) selected with a false alarm rate threshold of less than $1~\mathrm{yr}^{-1}$. This is identical to the set of observations used in KAGRA:2021duu. We consider only signals with BBH sources and the ambiguous GW190814 (LIGOScientific:2020zkf; LIGOScientific:2021usb). As GW190814’s secondary object may be above the expected maximum mass for a neutron star, we compare the results with and without GW190814, and investigate its potential as an outlier to the BBH population. GW200210_092254 (LIGOScientific:2021djp), which has a similar secondary mass and mass ratio to GW190814, is not included in our analysis as it does not satisfy the false alarm rate threshold.
We use a combined injection set of semi-analytic injections at the sensitivity of the first and second observing runs, and real injections at the third observing run sensitivity to account for selection effects (KAGRA:2021duu; GWTC3_sensitivity). We consider the real injections with a false alarm rate of less than $1~\mathrm{yr}^{-1}$ detectable, and the semi-analytic injections we cut in network signal-to-noise ratio of $10$ (LVKsensitivity2018).